OpenStudy (anonymous):

Evan is painting a stool that has a cylindrical seat and 4 rectangular prism legs. The seat has a diameter of 12 inches and a height of 2 inches. Each leg is 3 inches by 3 inches by 14 inches. How much area will Evan have to paint? Use 3.14 for pi. Show your work.

OpenStudy (anonymous):

@nathan917 please help me this is very important.

OpenStudy (nathan917):

1. surface area of the seat = 2 * 3.14 * 6^2 + 2 * 3.14*12 2. subtract 4 * 3*3 ( area of end of a leg) 3. add area of the 4 legs ( not including area of one end) = 4*( 3*3 + 14*3)

OpenStudy (nathan917):

Btw cool profile pic :P.

OpenStudy (anonymous):

469.44?

OpenStudy (nathan917):

One second making sure im right.

OpenStudy (nathan917):

1) Surface area of cylinder= 6^2 * 3.14 * 2 + 12 * 3.14 * 2 = 301.44 Subtract where legs touch seat --> 301.44 - 3 * 3 * 4 = 265.44 Surface area of legs = 3 * 3 * 14 * 4 = 504 Subtract where legs touch seat --> 504 - 3 * 3 * 4 = 468 265.44 + 468 = 733.44 sq. inches

OpenStudy (anonymous):

Oh, I failed then. ._. Lol

OpenStudy (nathan917):

No i should have explained it better :P. So i failed.

OpenStudy (anonymous):

We failed together :3

OpenStudy (nathan917):

Yes. So your in 7th Grade?.

OpenStudy (anonymous):

8th, I failed 7th grade math. How did you tell that by a question? lol

OpenStudy (nathan917):

Cause I'm in 7th. :P

OpenStudy (anonymous):

Ah. :p

OpenStudy (anonymous):

So this is what you do all day?

OpenStudy (nathan917):

No... :P